# Isotonic Layer: A Universal Framework for Generic Recommendation Debiasing

### v1

@@ Abstract.
Model calibration and debiasing are fundamental to ensuring the reliability and fairness of large-scale recommendation systems. While traditional methods like Platt Scaling (Platt, 1999) offer functional smoothness and Isotonic Regression (de Leeuw et al., 2009) provides non-parametric flexibility, both paradigms struggle to integrate into modern end-to-end deep learning pipelines: the former is often too restrictive for complex distributions, while the latter lacks the differentiability required for gradient-based optimization.
In this paper, we introduce the Isotonic Layer [^1] https://github.com/hailingc/Isotonic-Layer, a novel, differentiable framework that integrates piecewise linear fitting directly into neural architectures. By partitioning the feature space into discrete segments and optimizing non-negative slopes via a constrained dot-product mechanism, we instantiate a global monotonic inductive bias. Unlike standard neural layers, which lack inter-segment coordination, our approach enforces strict order properties, ensuring that model outputs remain logically consistent with critical features such as latent relevance, recency, or quality scores.
We further generalize this architecture by parameterizing segment-wise slopes as learnable embeddings. This enables the model to adaptively capture nuanced, context-specific distortions—such as position-based click-through rate (CTR) bias(Chen et al., 2021)—through specialized isotonic profiles. At its core, our approach utilizes a dual-task formulation that decouples the recommendation objective into (1) latent relevance estimation and (2) bias-aware calibration. By reframing debiasing as a differentiable score-calibration challenge, the Isotonic Layer serves as a functional bridge mapping biased observations back to their true underlying utility.
A major contribution of this work is the framework’s ability to perform highly granular, customized calibration. By learning embeddings for diverse context features—including display positions, device types, advertiser IDs, and objective types—the Isotonic Layer can achieve tailored calibration for arbitrary combinations of context features. This level of fine-grained control (e.g., per-advertiser calibration) is virtually impossible with traditional non-parametric methods. To address the inherent heterogeneity of modern systems, we extend this framework to Multi-Task Learning (MTL) environments with dedicated embeddings for distinct objectives. Extensive empirical evaluations on real-world datasets and large-scale production A/B tests demonstrate that the Isotonic Layer effectively mitigates systematic bias and enhances calibration fidelity without the risk of underfitting. Our approach significantly outperforms our production baseline in both predictive accuracy and ranking consistency.
@@ 1. Introduction
To the best of our knowledge, this work is the first to systematically integrate isotonic regression into the deep learning architecture by formulating it as a universal Isotonic Calibration Layer. By treating debiasing as an end-to-end calibration task, we present a generic framework that maintains structural monotonicity while leveraging the full representational power of deep ranking models.
@@ 1.4. Contributions
    Differentiable Isotonic Layer: We introduce a deep-learning-native Isotonic Layer that enforces global monotonicity within a neural framework. By conditioning the layer on high-dimensional context features (e.g., advertiser ID, device type) via learnable embeddings, we enable fine-grained, context-aware calibration that transcends the limitations of traditional non-parametric methods.
@@
    Efficient Piecewise Architecture: We implement the layer using a dot-product between a non-negative weight vector and a transformed input activation vector. This formulation leverages optimized BLAS operations, ensuring that the layer is computationally efficient for high-throughput production systems.
@@
    Handling Task Heterogeneity: To address the inherent heterogeneity of user behaviors, we scale the Isotonic Layer to Multi-Task Learning (MTL) environments through the introduction of task-specific isotonic embeddings. This architecture enables the model to adaptively adjust predicted score distributions by learning specialized monotonic profiles for distinct objectives. By doing so, the framework dynamically accounts for the unique distortion intensities associated with different engagement types — such as the high position-sensitivity of click-through rates (CTR) versus the intent-driven nature of conversion rates (CVR) — ensuring seamless and robust calibration across a unified MTL pipeline.
@@
    Dual-Task Formulation: We decouple recommendation into a latent relevance estimation task and a bias-aware calibration task, using the Isotonic Layer as a differentiable functional bridge and extended this formulation to a generic debiasing framework.
@@ 2. Related Work
Model calibration and debiasing are central challenges in recommender systems, where observed signals are systematically distorted by exposure, selection, and presentation effects. Prior work largely falls into three categories: post-hoc calibration, differentiable monotonic modeling, and causal or counterfactual debiasing.
2.1. Classical Calibration
@@
Post-hoc calibration methods such as Platt Scaling (Platt, 1999) apply a parametric logistic transformation to predicted scores. While efficient, these methods are too restrictive for the complex, non-linear distortions observed in modern recommender systems. Isotonic Regression (Zadrozny and Elkan, 2002) offers a non-parametric monotonic alternative, but relies on non-differentiable projection (e.g., PAVA), making it difficult to integrate into end-to-end training pipelines. In addition, isotonic regression often suffers from overfitting under data sparsity and produces staircase-shaped mappings that degrade ranking resolution.
2.2. Differentiable Monotonic Calibration
Recent work has explored embedding monotonic calibration into neural networks. Our previous paper, LiRank (et al., 2024) introduces a differentiable isotonic calibration layer by bucketizing logits and learning non-negative weights via backpropagation. Similarly, Deep Isotonic Promotion Network (DIPN) (et al., 2020) models monotonic incentive–response relationships using isotonic embeddings with additional smoothness regularization.
While effective, these approaches typically apply a single global monotonic function and focus on single-task settings. More recent neural monotonic calibrators based on unconstrained monotonic networks (Bai et al., 2025) further improve expressiveness, but remain limited to score-level calibration and do not address task heterogeneity or context-specific bias.
2.3. Debiasing via Causal and Counterfactual Learning
A parallel line of work addresses debiasing through causal inference and counterfactual learning. Classical approaches include Position-as-Feature (PAF) modeling and Inverse Propensity Scoring (IPS) (et al., 2017), which reweight training samples to simulate randomized exposure. However, PaF suffers from inference-time bias shift, while IPS-based methods are high-variance and sensitive to propensity estimation errors.
Recent deep learning approaches improve robustness by modeling bias explicitly through representation learning or meta optimization. These include disentangled or variational models that separate preference from bias factors (Guo et al., 2023), adversarial or propensity-learning frameworks (Zhu et al., 2025), and general debiasing meta-frameworks such as AutoDebias (Chen et al., 2021), which jointly optimize reweighting, imputation, and prediction. While powerful, these methods primarily operate on the data distribution or latent representations and do not impose explicit structural constraints on the score transformation itself.
Beyond monotonic calibration and counterfactual debiasing, recent work has explored multifactorial bias correction in recommendation, where both popularity and positivity biases are jointly modeled to improve fairness and robustness (Huang et al., 2024a; Mansoury et al., 2026). Contrastive representation learning has been applied to mitigate dual biases such as popularity and conformity (Huang et al., 2024b), while invariant learning frameworks aim to disentangle invariant user preferences from biased interactions using imputation and knowledge distillation (Daniil et al., 2025). Additionally, thorough analyses and surveys on popularity bias shed light on the broader impacts of algorithmic bias in recommender systems and their mitigation strategies (Carnovalini et al., 2025; Klimashevskaia et al., 2024).
2.4. Multi-Task Learning and Bias Heterogeneity
@@
Modern industrial recommenders are inherently multi-task, optimizing signals with vastly different bias profiles (e.g., clicks vs. conversions). Prior work in multi-task debiasing focuses on mitigating negative transfer via task decoupling or causal regularization (et al., 2025). However, most approaches implicitly assume a shared calibration or bias correction mechanism across tasks, despite large differences in exposure sensitivity.
2.5. Summary
@@
In contrast to prior work, our approach treats debiasing as a differentiable monotonic calibration problem and embeds it directly into multi-task neural architectures. By learning context- and task-conditioned isotonic embeddings, the proposed Isotonic Layer enables fine-grained, per-objective calibration within a unified multi-task framework—addressing bias heterogeneity that global monotonic or causal reweighting methods do not explicitly handle.
@@ 3. Isotonic Layer Architecture Overview
We propose an Isotonic Layer, a differentiable neural network module that enforces monotonically non-decreasing outputs with respect to a scalar input, illustrated in Figure 1(b). Unlike classical isotonic regression methods that rely on exact projection operators, the proposed layer adopts a bucket-based cumulative construction that is fully compatible with end-to-end gradient-based optimization.
The layer discretizes the input domain into fixed-width buckets and associates each bucket with a learnable non-negative weight. Monotonicity is guaranteed by (i) cumulative activation of buckets as the input increases and (ii) explicit non-negativity constraints on bucket weights enforced via a ReLU parameterization.
@@ 3.1. Mathematical Formulation
In practice, the Isotonic Layer can be trained on the upstream model’s dataset or a separate holdout set. While joint training optimizes relevance and calibration simultaneously, using a distinct dataset decouples structural bias from relevance learning. This separation is particularly effective for correcting systematic distortions—such as position or presentation bias—rather than merely improving raw accuracy. Consistent with prior work in post-hoc calibration and counterfactual learning, isolating calibration from estimation enhances robustness and generalization under distribution shift.
@@ 3.2. Monotonicity Guarantee
For any two inputs $x_{1}\leq x_{2}$, their corresponding activation vectors satisfy
(9) $\mathbf{a}(x_{1})\leq\mathbf{a}(x_{2})\quad\text{(element-wise)}.$
Since all bucket weights are non-negative ($w^{+}_{j}\geq 0$), the pre-activation output is monotonic:
(10) $z(x_{1})\leq z(x_{2}).$
Because the sigmoid function is strictly monotonic, the final output preserves this ordering:
(11) $y(x_{1})\leq y(x_{2}).$
Thus, monotonic non-decreasing behavior is guaranteed by construction.
@@ 3.3. Expressiveness and Empirical Illustration
To evaluate the expressive capacity of the Isotonic Layer, we first conduct a synthetic experiment using the monotonic target function $y=x^{2}$ over $x\in[0,1]$. To mimic the score distribution commonly observed in recommendation systems, the input is transformed into logit space via $\mathrm{logit}(x)=\ln\frac{x}{1-x}$ prior to being passed into the layer. The model is optimized using binary cross-entropy loss. As shown in Figure 2, the Isotonic Layer achieves a near-perfect approximation of the quadratic target, demonstrating that the proposed piecewise linear construction retains strong expressive power despite monotonic constraints.
Figure 2. Expressive capacity of the Isotonic Layer on a monotonic synthetic task ($y=x^{2}$).
Demo of monotonic task
We further evaluate robustness under non-monotonic distortions by constructing a piecewise target function defined as $y=x^{2}$ for $x\leq 0.95$ and $y=(1.9-x)^{2}$ for $x>0.95$. This setup simulates overfitting artifacts that may arise in real-world models due to noise or data sparsity. As illustrated in Figure 3, the Isotonic Layer effectively suppresses local inversions while recovering the optimal monotonic approximation, highlighting its ability to provide structural regularization under noisy or inconsistent inputs.
@@
Unlike standard multi-layer perceptrons, whose parameters operate locally and independently, the Isotonic Layer imposes a global structural constraint. Because the activation at input $x$ accumulates contributions from all preceding buckets, the output depends jointly on the entire prefix of the weight embedding. This induces strong spatial coupling across segments.
As a result, the model cannot arbitrarily modify the function shape in isolated regions, even under data sparsity. This global dependency acts as an implicit regularizer, preventing local overfitting or monotonicity violations while retaining high expressive capacity through fine-grained bucketization. Empirically, this property allows the Isotonic Layer to generalize reliably in regions with limited observations, a common challenge in large-scale recommendation systems.
@@ 3.5. Context-Conditioned Isotonic Embeddings
A key advantage of the Isotonic Layer is that its bucket weights function as a learnable embedding parameterizing a monotonic transformation. This allows for context-aware calibration, where the isotonic function adapts dynamically based on auxiliary features like display position or device type (see Figure 1(c)). For a context variable $c$, we define a context-conditioned weight embedding:
(12) $\mathbf{w}(c)=\mathbf{E}(c)\in\mathbb{R}^{N},$
where $\mathbf{E}(\cdot)$ is implemented via an embedding lookup or a lightweight neural projection. Each context induces a distinct isotonic calibration curve while preserving global monotonicity through the non-negativity constraint:
(13) $\mathbf{w}^{+}(c)=\mathrm{ReLU}(\mathbf{E}(c)).$
The calibrated output is then computed as
(14) $y=\mathrm{IsotonicLayer}(x\,;\,\mathbf{w}^{+}(c),b(c)),$
where both the bucket weights and bias may be conditioned on context. (code implementation see appendix A)
Example: Position Bias.
In ranking systems, display position $p$ often introduces observational bias. By associating each position with a learnable isotonic embedding $\mathbf{E}_{p}$, the model captures position-specific monotonic distortions of the latent relevance score. This mechanism enables the calibration layer to dynamically ”stretch” or ”compress” score distributions, effectively modeling position bias as a learned monotonic transformation:
(15) $\hat{y}=f_{iso}(r;\mathbf{E}_{p})$
where $r$ is the raw relevance score and $f_{iso}$ is the position-conditioned isotonic function.
@@ 3.6. Dual-Tower Debiasing Architecture
We operationalize bias-aware learning by decomposing the model into two complementary components connected via the Isotonic Layer, as illustrated in Figure 1(d):
  • Relevance Tower. A standard neural network that estimates the latent utility $r$ of an item given user and item features, independent of exposure bias.
  • Isotonic Calibration Layer. A context-conditioned monotonic mapping that transforms latent utility into the observed interaction space, capturing systematic distortions such as position bias or interface effects.
Formally, the observed prediction is given by
(16) $\hat{y}=f_{iso}(r\,;\,c),$
where the context $c$ controls the isotonic embedding. This decoupling allows the model to learn bias patterns explicitly while preserving a clean separation between relevance estimation and calibration.
Bias Neutralization at Inference.
A key practical advantage of this formulation is that bias can be neutralized at inference time. By fixing the context embedding to a reference value (e.g., a canonical display position) or by bypassing the calibration layer entirely, the system can produce rankings driven purely by latent utility:
(17) $\hat{r}\approx r.$
This enables counterfactual evaluation, fair ranking, and policy simulation without retraining the relevance model, making the approach particularly attractive for large-scale production systems.

### v2

# Isotonic Layer: A Unified Framework for Recommendation Calibration and Debiasing

@@ Abstract.
Model calibration and debiasing are fundamental yet operationally expensive challenges in large-scale recommendation systems. Existing approaches treat them as separate problems requiring distinct infrastructure: post-hoc calibration pipelines, propensity estimation workflows, and per-segment model farms. We introduce the Isotonic Layer [^1] https://github.com/hailingc/Isotonic-Layer, a differentiable piecewise-linear module that unifies both problems within a single, lightweight architectural component—requiring no additional data preprocessing, no propensity estimation, and no separate calibration pipelines.
The core insight is elegant: by parameterizing non-negative bucket weights as learnable context embeddings, the model automatically learns all calibration and debiasing functions end-to-end from standard training data. Swapping in a different embedding (position, device type, advertiser ID, or any combination) instantly yields calibration tailored to that sub-segment—at arbitrary granularity in any high-dimensional feature space—with no engineering changes beyond a single embedding lookup. The same layer handles post-hoc calibration, position debiasing, and heterogeneous multi-task bias correction within one unified framework. This paper does not claim state-of-the-art over every benchmark; it offers a principled, practical simplification: a plug-and-play solution that replaces fragmented, high-maintenance calibration infrastructure with a single end-to-end trainable component. Extensive production A/B tests confirm significant improvements in predictive accuracy, calibration fidelity, and ranking consistency.
@@ 1. Introduction
The Isotonic Layer was originally conceived by the first authors of this paper and first deployed as a component within LiRank (et al., 2024), LinkedIn’s production ranking system. Due to scope constraints in that system paper, only a global, single-task variant of the layer was described, without formal analysis or generalization to context-conditioned embeddings, multi-task settings, or a unified calibration and debiasing framework. This paper presents the complete formalization of the Isotonic Layer—encompassing its theoretical foundations, monotonicity guarantees, context-conditioned extensions, and dual-tower debiasing formulation—as a unified end-to-end calibration framework that maintains structural monotonicity while leveraging the full representational power of deep ranking models.
@@ 1.3. Proposed Solution: The Isotonic Layer
1.4. Design Philosophy
The central value of this work is unification through simplicity. Calibration and debiasing are not treated as separate problems requiring separate infrastructure—they are both instances of the same learned monotonic transformation, automatically derived from training data without any manual data engineering. Specifically:
  • No data preprocessing required. Unlike IPS-based methods that need propensity estimation or randomized logging, the Isotonic Layer learns bias patterns directly from observational data.
  • Arbitrary sub-segment calibration via a single embedding change. Calibrating for a new context (e.g., a specific advertiser, device, or position) requires only adding an embedding lookup—no new model, no new pipeline.
  • Fully automatic, end-to-end. All calibration and debiasing functions are learned by the model itself during standard training. No post-hoc fitting, no manual curve construction.
  • Plug-and-play for any multi-task setup. Task-specific embeddings handle heterogeneous bias profiles across arbitrary objectives within one unified architecture.
@@ 1.5. Contributions
    Unified Calibration & Debiasing Framework: A single differentiable layer replaces the need for separate calibration pipelines, propensity models, and per-task sub-models, unifying both problems under one end-to-end trainable component.
@@
    Context-Conditioned Isotonic Embeddings: Bucket weights conditioned on arbitrary context features (e.g., advertiser ID, device type, display position) enable fine-grained per-segment calibration at any granularity in high-dimensional spaces—impossible with classical methods.
@@
    Handling Task Heterogeneity: Task-specific isotonic embeddings learn heterogeneous bias profiles for distinct objectives (CTR vs. CVR) within a unified MTL pipeline, dynamically accounting for the unique distortion intensities of each engagement type.
@@
    Dual-Task Formulation: We decouple recommendation into latent relevance estimation and bias-aware calibration, using the Isotonic Layer as a differentiable functional bridge, with bias fully neutralizable at inference without retraining.
@@ 2. Related Work
Classical calibration. Platt Scaling (Platt, 1999) applies a parametric logistic transformation to predicted scores but is too restrictive for non-linear distortions in modern recommenders. Isotonic Regression (Zadrozny and Elkan, 2002) offers a non-parametric monotonic alternative but relies on non-differentiable projection (PAVA), resists end-to-end training, and produces staircase mappings that degrade ranking resolution.
Differentiable monotonic calibration. The Isotonic Layer was originally conceived by its first authors (Cheng and Yang) and deployed within LiRank (et al., 2024) (patent US20250278623A1) as a single-task, global instantiation; this paper provides its complete formal treatment and generalization. DIPN (et al., 2020) similarly uses isotonic embeddings for monotonic incentive–response modeling. Both apply a single global function and target single-task settings. Unconstrained monotonic networks (Bai et al., 2025) improve expressiveness but remain limited to score-level calibration without addressing task heterogeneity or context-specific bias.
Causal and counterfactual debiasing. Position-as-Feature (PAF) and Inverse Propensity Scoring (IPS) (et al., 2017) reweight samples to simulate randomized exposure, but suffer from inference-time distribution shift and high propensity estimation variance, respectively. Recent approaches model bias through disentangled representations (Guo et al., 2023), adversarial learning (Zhu et al., 2025), or meta-optimization (AutoDebias (Chen et al., 2021)), and address multifactorial biases (Huang et al., 2024a; Mansoury et al., 2026; Huang et al., 2024b; Daniil et al., 2025; Carnovalini et al., 2025; Klimashevskaia et al., 2024). These methods operate on data distributions or latent representations without imposing structural constraints on score transformations.
@@
Multi-task debiasing. Prior work mitigates negative transfer via task decoupling or causal regularization (et al., 2025), but most methods assume a shared calibration mechanism across tasks despite large differences in exposure sensitivity across objectives such as clicks versus conversions.
@@
Our approach treats calibration and debiasing as the same differentiable monotonic transformation, unified in a single plug-and-play layer requiring no preprocessing or propensity estimation. Context- and task-conditioned isotonic embeddings enable fine-grained per-objective calibration within a unified multi-task framework—addressing bias heterogeneity that global monotonic or causal reweighting methods do not handle.
@@ 3. Isotonic Layer Architecture Overview
3.1. Two Applications of the Same Layer
The Isotonic Layer is a single architectural component that serves two distinct applications in recommendation systems: calibration (Section 4) and debiasing (Section 5). In both cases, the mechanism is the same—an embedding lookup on a context feature parameterizes the isotonic bucket weight vector, and the model learns the corresponding monotonic transformation end-to-end. The important difference lies in how the layer is deployed across training and serving:
  • Calibration: the Isotonic Layer takes calibration context features (e.g., platform type, user segment) as input and is active during both offline training and online serving. The same context-conditioned calibration function learned during training is applied identically at inference, continuously correcting systematic score distortions without any additional pipeline.
  • Debiasing: the Isotonic Layer takes bias-inducing context features (e.g., feed position, device type) as input and forms a dedicated Isotonic Calibration Head trained alongside a standard Inference Head. It is active only during offline training, where it provides a corrective learning signal that shapes the shared representation. At online serving time, the Isotonic Head and its debiasing features are removed entirely; only the Inference Head is deployed, producing rankings driven by pure latent relevance free of positional or device bias.
The train/serve asymmetry reflects the fundamental difference: calibration distortions persist at serving time and must be corrected there, whereas debiasing signals (position, device) are logging-policy artifacts that must not influence online predictions. In both cases the underlying building block is identical, requiring no additional pipeline, preprocessing, or propensity estimation.
We propose an Isotonic Layer (Figure 1(b)): a differentiable, bucket-based neural module that enforces monotonically non-decreasing outputs via cumulative bucket activations and ReLU-constrained non-negative weights, fully compatible with end-to-end gradient-based optimization.
@@ 3.2. Mathematical Formulation
The layer can be trained jointly with the upstream model or on a separate holdout set; the latter decouples calibration from relevance learning, improving robustness under distribution shift.
@@ 3.3. Monotonicity Guarantee
For any $x_{1}\leq x_{2}$, the activation vectors satisfy $\mathbf{a}(x_{1})\leq\mathbf{a}(x_{2})$ element-wise by construction. Since $w^{+}_{j}=\mathrm{ReLU}(w_{j})\geq 0$, the dot product preserves this ordering: $z(x_{1})\leq z(x_{2})$, and the strictly monotonic sigmoid yields $y(x_{1})\leq y(x_{2})$. Monotonicity is therefore guaranteed by construction—it is precisely the ReLU on bucket weights that enforces non-negativity and thereby isotonicity.
Generalization: Unconstrained Piecewise Linear Fitting.
The ReLU constraint is not always necessary. In applications where a monotonicity guarantee is not required—for example, fitting a general score transformation or modeling non-monotonic feature interactions—the ReLU can simply be removed, allowing bucket weights $w_{j}$ to take arbitrary real values. The layer then becomes a universal piecewise linear function approximator, capable of fitting any continuous shape over the input domain. This makes the same architecture directly applicable as a general-purpose learned transformation beyond the calibration and debiasing settings described in this paper.
@@ 3.4. Expressiveness and Empirical Illustration
Figure 2 shows the layer achieving near-perfect approximation of $y=x^{2}$ (inputs in logit space, optimized with BCE), confirming strong expressive power under monotonic constraints. Figure LABEL:fig:isotonic2 evaluates robustness to non-monotonic distortions via a piecewise target ($y=x^{2}$ for $x\leq 0.95$; $y=(1.9-x)^{2}$ for $x>0.95$): the layer suppresses local inversions while recovering the optimal monotonic fit, demonstrating structural regularization under noisy inputs.
@@
Unlike MLPs whose parameters act locally, the Isotonic Layer imposes a global structural constraint: the output at $x$ depends on the cumulative prefix of all preceding bucket weights, inducing spatial coupling across segments. This acts as an implicit regularizer—preventing local overfitting or monotonicity violations even under data sparsity—while retaining high expressive capacity through fine-grained bucketization.
@@ 3.6. Context-Conditioned Isotonic Embeddings
The bucket weights serve as a learnable embedding parameterizing a monotonic transformation (Figure 1(c)). For a context variable $c$, an embedding lookup or lightweight projection yields $\mathbf{w}^{+}(c)=\mathrm{ReLU}(\mathbf{E}(c))\in\mathbb{R}^{N}$, giving a distinct isotonic calibration curve per context while preserving monotonicity:
(9) $y=\mathrm{IsotonicLayer}(x\,;\,\mathbf{w}^{+}(c),\,b(c)).$
For example, associating each display position $p$ with embedding $\mathbf{E}_{p}$ captures position-specific score distortions as $\hat{y}=f_{\mathrm{iso}}(r;\,\mathbf{E}_{p})$, where $r$ is the raw relevance score. (See Appendix C for implementation details.)
@@ 3.7. Dual-Tower Debiasing Architecture
As illustrated in Figure 1(d), the model decomposes into a Relevance Tower estimating latent utility $r$ and an Isotonic Calibration Layer mapping $r$ to the observed interaction space via context $c$:
(10) $\hat{y}=f_{\mathrm{iso}}(r\,;\,c).$
At inference, bypassing the calibration layer yields rankings driven purely by latent utility ($\hat{r}\approx r$), enabling counterfactual evaluation and fair ranking without retraining.